OpenStudy (anonymous):

Please help! Thanks! z1=-3+4i z2=2-i Perform the indicated operations and write the answer in the form a + bi. z1/z2= ?

OpenStudy (anonymous):

do you know what conjugate means? for example, what's the conjugate of a-b ?

OpenStudy (mertsj):

\[\frac{-3+4i}{2-i}\times \frac{2+i}{2+i}\]

OpenStudy (anonymous):

Thanks! Can you please tell me what -15/2-5i/2 is simplified?